Fermi National Laboratory


Accelerator Update

Monday June 21
The day shift began with Operations monitoring store 3586 and the Recycler stash, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

At 9:54 AM, MiniBooNE made the first of what would become many accesses for a special study.

A Main Injector (MI) expert began a one-hour MECAR study at 10:16 AM.

Booster experts investigated the regulation problem with a power supply (B:SEXL4). They reported that the repair wouldn't be easy. The supply was left off.

Operations reestablished beam to MiniBooNE at 12:18 PM and to the Meson MTest experiment at 1:11 PM.

Tevatron (TeV) experts reported at 1:34 PM that they had made changes to the collimator movement sequencer.

At 2:02 PM, the Collider Detector at Fermilab (CDF) and D-Zero (D0) both reported a collider beam problem. Experts thought at first it was due to a separator spark, but later discovered it was due to a heavy truck driving by CDF on the Ring Road.

At 4:32 PM, Recycler experts began an instability study. This study required there to be no ramp or beam in MI.

The Recycler study finished and Pbar resumed stacking about 5:50 PM.

Operations established beam to Meson MCenter experiment at 8 PM

A Pbar expert began a stacking study at 10:15 PM. This interrupted stacking and also halted all Switchyard beam.

Pbar resumed stacking at 11:37 PM.

Tuesday June 22
The midnight shift began with Operations monitoring store 3586, with Pbar stacking, MiniBooNE, MTest, and MCenter taking beam.

At 5:16 AM, A Booster RF station (BRF1) tripped off and wouldn't reset. Operators bypassed the station and called in an expert who determined at 6:09 AM that the repair would require an access. The Run Coordinator determined that this repair would be made later.

Operations terminated store 3586 at 8:16 AM and put the TeV into shot setup at 9:11 AM.

Operations established store 3588 at 11:36 AM with an initial luminosity of 84.78E30. This set A NEW LUMINOSITY RECORD. Pbar resumed stacking.

Operations reestablished beam to MTest at 12:05 PM.

Wednesday June 23
The midnight shift began with Operations monitoring store 3688 and the stash, with Pbar stacking, and with MiniBooNE and MTest taking beam.

At 1:35 AM, Operations reported that a Recycler corrector power supply (R:CE401) tripped off. Eventually this caused the loss of the stash.

Plans for Today
The plans for today are to terminate the store at 5 PM and then allow a study period.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/23/2004   email Fermilab